The directions say find an equation in standard form of the line containing point p and having slope m. The problem is P(2,1), m=-1
Answer by oberobic(2304)

m = -1 is the slope.
so,
y = -x + b
.
You can use the point given to find 'b'.
.
1 = -2 + b
b = 3
.
y = -x + 3
.
That is slope-intercept form.
.
Add x to both sides to get:
.
x + y = 3
.
That is "standard" form of a linear equation.
.
In general terms, the "standard form" is: Ax + By = C, where A, B, and C are constants.
.
Done.